OpenStudy (anonymous):

The two rectangles are similar. Which is a correct proportion for corresponding sides?

OpenStudy (anonymous):

D) GD/MJ = DE/JK B) GD/MJ = EF/ML C) GD/MJ = DE/ML D) GD/MJ = EF/JK